Cameron has saved $18,000 and plans to invest it in two accounts: a savings account that pays 5.2% annual interest and a stock that promises a 9% annual return. If Cameron earns $1,430 in interest after one ear, how much did she invest in each account?
Answer by mananth(16949)

Investment I 5.20% per annum
Investment II 9.00% per annum

x + y 18000 ------------------------1
5.20% x + 9.00% y = $1,430.00
Multiply by 100
5.2 x + 9 y = $143,000.00 --------2

Multiply (1) by -5.2
we get

-5.2 x -5.2 y = -93600.00

Add this to (2)

0 x 3.8 y = $49,400.00

divide by 3.8

y = $13,000.00 investment at 9.00%
Balance $5,000.00 investment at 5.20%
CHECK
$5,000.00-------$260.00
$13,000.00------$1,170.00
Total--------- $1,430.00